OpenStudy (anonymous):

log5 (x+2) - log5 10= log5 100

OpenStudy (anonymous):

log5 (x + 2) = log5 100 + log5 10 log5 (x + 2) = log5 1000 x + 2 = 1000 x = 1000 - 2 = 998